# Do ImageNet-trained models learn shortcuts? The impact of frequency shortcuts on generalization

### v1

@@ 5 Conclusions
We proposed the first method for analysis of frequency shortcuts, HFSS, that enables the inspection of shortcut learning in large-scale models and datasets with thousands classes. HFSS is more time-efficient and effective in finding frequency shortcuts compared to existing approaches. We investigate frequency shortcut learning in models trained on ImageNet and relate their results to robustness and generalization performance under different OOD conditions. The impact of frequency shortcuts on model generalization depends on the specific OOD scenarios. Existing models yield good performance on generalization benchmarks when texture information, corresponding to most frequency shortcuts, is mostly preserved in OOD data. Instead, frequency shortcuts impair the generalizability of models to images with rendition changes. This highlights the limitation of current OOD performance evaluation benchmarks, which need to explicitly take into account the impact that frequency shortcuts have on model performance. HFSS provides a tool to bridge this gap and extend the rigor of model generalization evaluation.

### v2

@@ 5 Conclusions
We proposed the first method for analysis of frequency shortcuts, HFSS, that enables the inspection of shortcut learning in large-scale models and datasets with thousands of classes. HFSS is more time-efficient and effective in finding frequency shortcuts compared to existing approaches. We investigate frequency shortcut learning in models trained on ImageNet and relate their results to robustness and generalization performance under different OOD conditions. The impact of frequency shortcuts on model generalization depends on the specific OOD scenarios. Existing models yield good performance on generalization benchmarks when texture information, corresponding to most frequency shortcuts, is mostly preserved in OOD data. Instead, frequency shortcuts impair the generalizability of models to images with rendition changes. This highlights the limitation of current OOD performance evaluation benchmarks, which need to explicitly take into account the impact that frequency shortcuts have on model performance. HFSS provides a tool to bridge this gap and extends the rigor of model generalization evaluation.
Supplementary Material
@@ C.1 IN-1k
Number of classes per stage.
Tab. 5 presents the number of shortcut classes at each stage. As the threshold $t$ increases, the count of shortcut classes declines. Notably, at stage 6 with $t=0.9$, the number drops to 1-3, indicating that such strong shortcuts are uncommon.
Table 5: The number of shortcut classes per stage in IN-1k.
Stage $t$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
ResNet18
3 993 950 894 792 642 516 317 175 40
4 922 826 701 539 381 276 133 62 9
5 647 445 301 203 111 61 23 12 3
6 343 196 116 60 22 16 7 3 1
ResNet50
3 997 990 956 904 782 673 469 305 82
4 976 903 792 639 457 339 175 95 12
5 809 604 445 308 174 110 45 18 2
6 421 253 149 78 33 18 5 1 1
CCT
3 1000 997 989 958 890 820 648 475 166
4 998 985 948 885 755 636 421 256 72
5 973 906 778 622 442 336 152 82 15
6 781 566 408 277 142 88 34 15 3
ViT-b
3 1000 989 977 930 820 720 518 343 101
4 987 962 900 772 633 491 297 171 33
5 861 674 514 368 225 149 70 25 6
6 520 317 183 110 54 30 7 1 1